OpenStudy (haltiamali):

[h o g](x) h(x)=x^2 + 2x - 1 g(x) = x - 2 Help please?? :/

OpenStudy (haseeb96):

[h o g ] (x) = h ( x-2 ) now hog means put the value or equation of g into the equation of h on the place of x okay? so the equation would be [hog](x)=(x-2)^2 +2x-1 now expand the equation and show me

OpenStudy (haltiamali):

so then it would be (x-2)^2 + 2(x-2)-1?

OpenStudy (haltiamali):

or do you only put the g in for the first x?

OpenStudy (anonymous):

yeh

OpenStudy (haltiamali):

then do I distribute?

OpenStudy (haseeb96):

no sorry my mistake it would put in all the place of x [hog] (x)= (x-1)^2 +2(x-2)-1

OpenStudy (haseeb96):

my mom called me so i forgot that sorry that was my mistake

OpenStudy (haltiamali):

the finally answer is x^2+2x-9, right?

OpenStudy (haseeb96):

now expand it

OpenStudy (anonymous):

but why u put x-1 in the place of x^2 @Haseeb96

OpenStudy (haseeb96):

(a-b)^2 = a^2 -2ab +b^2 use this formula for that power 2 m

OpenStudy (haseeb96):

sorry again mistake there would be x-2

OpenStudy (haseeb96):

[hog](x) = (x-2)^2 +2(x-2) -1

OpenStudy (haseeb96):

now expand it

OpenStudy (haseeb96):

no i was confused because i was helping other

OpenStudy (haltiamali):

I got x^2 - 4 + 2x - 4 - 1 after that and reduced it to x^2 + 2x - 9

OpenStudy (haltiamali):

right? :P

OpenStudy (haseeb96):

[hog](x)= (x^2 - 4x + 4) + (2x -4) -1 = x^2 -2x -1 this is correct answer

OpenStudy (haseeb96):

@rosebird please tag me there